OpenStudy (anonymous):

Satellite73, could you please help me?youre the only one thats good at this precalc stuff lol

OpenStudy (anonymous):

Or anyone thats good at it, please? :(

OpenStudy (anonymous):

what do you need help with.

OpenStudy (anonymous):

specifically

OpenStudy (anonymous):

Use the exponential model P=54.7e ^.19t to determine when P=123.68. Then convert to base b

OpenStudy (anonymous):

that is easy 123.68/54.7 = e^(.19t) = ln(123.68/54.7) = .19t t = ln(123.68/54.7)/.19

OpenStudy (anonymous):

stuff with exponential growth and decay models.

OpenStudy (anonymous):

I posted somme other ones that people didnt know how to answer:?can you look at them?

OpenStudy (anonymous):

I guess..

OpenStudy (anonymous):

can you just post the questions in this thread you need help with specifically

OpenStudy (anonymous):

In 1990, the population of africa was 643 million and by 2000 it had grown to 813 million. Write an exponential model in base b for the population of africa t after 1990. What is the annual percent increase?

OpenStudy (anonymous):

so we use the exponential model 813=643e^(10*r) we are trying to find the rate so do what I did before with this equation 813=643e^(10*r) 813/643 = e^(10r) ln(813/643) = 10r ln(813/643)/10 = r here are some lecture notes that cover exponential functions and the type of questions that you are covering

OpenStudy (anonymous):

yeah i have a bunch of notes, i jst dont understand it

OpenStudy (anonymous):

When you read the question just right down all the variables that it gives you you know the intiial population is n = 643 Principle = 813 Time = 10 years (2000 - 1990) and you should know the formula P(t) = ne^(tr)

OpenStudy (anonymous):

what are you not getting?

OpenStudy (anonymous):

i just did the calculations for that one and i got .02345, and that doesnt seem right at all

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

the africa population one

OpenStudy (anonymous):

its right, I just checked

OpenStudy (anonymous):

the population is increasing exponentially by 2%.

OpenStudy (anonymous):

then what is the model for it?

OpenStudy (anonymous):

You can check by inserting .023 into r in the equation 643e^(10*0.023) = 813

OpenStudy (anonymous):

You are modeling for population growth, so you are finding at what percent exponentially the population increased, with that rate you can insert 20 for t and know how people in africa starting with 643million will be in africa after 20 years

OpenStudy (anonymous):

but it doesnt make sence that its only 2 percent when it went from 643, to 813, that seems like a higher percentage for everyyear increase

OpenStudy (anonymous):

yes it does it is an exponential function. Look at the graph e^(x)

OpenStudy (anonymous):

5^(5) = (5)(5)(5)(5)(5) = (25)(5)(5)(5) 125(5)(5) 625(5) = 3125 see how large this number got

OpenStudy (anonymous):

well then the second part of that question is write an exponential model in base e for the popul;ation of afrca t years after 1990 . what is the continuous annual increse?

OpenStudy (anonymous):

Just multiply .023 by 100 to convert it to percent, thus 2.3 percent increase each year

OpenStudy (anonymous):

butfor continuously, its a different equation

OpenStudy (anonymous):

is that the full question?

OpenStudy (anonymous):

ehi dont even know, i have a feeling were getting all this wrong

OpenStudy (anonymous):

No the last two questions are right but you don't have to believe me

OpenStudy (anonymous):

i dont know, continuously is a different equation though let me find it

OpenStudy (anonymous):

its the pe^rt formula

OpenStudy (anonymous):

that is the formula I used before...

OpenStudy (anonymous):

only I formated it P=ne^(rt)

OpenStudy (anonymous):

but thats only for continuouslyt

OpenStudy (anonymous):

Are they just tell you to state the function? because I already did that 643e^(t*0.023) =P(t)

OpenStudy (anonymous):

If you graphed that function it would show you how the african population increases over the years after 1990

OpenStudy (anonymous):

it tells me to put the model, and the percent increase on both base e and base b

OpenStudy (anonymous):

can I see the question?